# Language Modeling Is Compression

### v1

@@ 3.5 Sequential Evolution of In-Context Compression
Language models take a very different “approach” to compression compared to classical compressors. Classical compressors have a small program size and optimize for a large context length to exploit sequential dependencies in the data. In contrast, foundation models consist of billions of parameters, which enable rapid adaptation in their (relatively) short context window (Genewein et al., 2023). Thus, arithmetic coding-based compressors rely heavily on the predictive models’ in-context learning capabilities to achieve competitive compression performance. We investigate this phenomenon in Fig. 6, which visualizes the compression rate across sequence lengths for gzip, Chinchilla 1B and a Transformer pretrained on enwik8. Intuitively, the longer the sequence, the more data the model can process in its context, and therefore, the better the compression. As expected, most compression rates decrease quickly with increasing sequence length, indicating that the models learn some data statistics in-context, without any gradient-based training. As in Table 1, the Chinchilla model achieves the best compression rates accross all three data modalities and sequence lengths.
@@ 3.6 Tokenization Is Compression
Table 2: Raw compression rates (compressed size / raw size) on enwik9 for Transformers trained on enwik8 with different tokenizers, ASCII and byte-pair encoding (BPE), with various vocabulary sizes. Transformers compress better with simpler tokenizers. However, larger vocabulary sizes reduce the length of the sequence more, meaning more information can be packed into the context.
Raw Compression Rate (%)
Tokenization 200K 6.4M 38M
ASCII 22.9 13.6 6.4
BPE 1000 25.4 14.8 6.9
BPE 2000 25.6 15.7 7.4
BPE 5000 23.1 17.1 8.7
BPE 10000 21.3 17.0 8.9
BPE 20000 19.3 16.4 9.0
@@
Increasing the number of tokens (i.e., the “alphabet size”) reduces the length of the sequence and thus increases the amount of information in a models context. However, decreasing the sequence length comes at a price: the number of tokens is larger, which makes the prediction task more challenging since reducing the entropy of the conditional distribution $\rho(x_{i}\mid x_{<i})$ is increasingly difficult for larger alphabet size. In theory, as the tokenization is a lossless compression, the two effects should compensate. In practice, we observe that if the model is small, increasing the number of possible tokens boosts the compression performance. In contrast, for bigger models, it seems that the converse happens: having a larger token vocabulary harms the final compression rate of the model. Nevertheless, short sequence lengths also help Transformers since their time complexity scales quadratically with context length, and it has been shown they do not generalize well to long contexts (Delétang et al., 2023; Ruoss et al., 2023). This explains why most practical Transformer implementations still use some form of tokenization, e.g., SentencePiece (Kudo & Richardson, 2018).
@@ 4 Related work
Leveraging Shannon’s source coding theorem (Shannon, 1948), a plethora of approaches exploit the connection between prediction and compression. For example, context-tree weighting (CTW) (Willems et al., 1995) mixes the predictions of many underlying Markov models to achieve lossless compression via arithmetic coding (Pasco, 1977; Rissanen, 1976). Similarly, prediction by partial matching (PPM) (Cleary & Witten, 1984) also leverages arithmetic coding, but uses a contiguous context matching method to create probability distributions based on the history of characters in a sequence. Likewise, PAQ8 (Knoll & de Freitas, 2012) uses a weighted combination of predictions from a large number of models (most of them based on context matching, but unlike PPM also noncontiguous context matches). In a different setting, Veness et al. (2015) demonstrated how to employ compression to obtain value estimates of a policy in an environment. Frank et al. (2000) and later Teahan & Harper (2003) introduced the idea of classification with compressors. Recently, Jiang et al. (2023) applied this technique with NLP tasks, paired with a k-nearest-neighbour algorithm. The results are surprisingly good for simple general purpose compressors like gzip. Jiang et al. (2022) exploit the same idea but train the compressor on a vast amount of unlabeled data first. Finally, van den Oord & Schrauwen (2014) apply arithmetic coding to image compression using Student distribution mixtures and Gaussian processes as predictors.
@@
Prior work demonstrated that neural predictive distributions can be employed to perform lossless compression via arithmetic coding (Schmidhuber & Heil, 1996; Mahoney, 2000; Knoll, 2014; Cox, 2016; Schiopu et al., 2018; Goyal et al., 2019; Liu et al., 2019; Mentzer et al., 2019, 2020; Schiopu & Munteanu, 2020; Rhee et al., 2022; Mikolov, 2012). Similarly, neural networks were also shown to achieve strong lossless compression rates when replacing arithmetic coding with asymmetric numeral systems (Hoogeboom et al., 2019; Kingma et al., 2019; Townsend et al., 2019; Barzen et al., 2022). While these approaches assume the existence of a separate training set, a different line of work investigated arithmetic coding-based neural compression in a purely online fashion, i.e., training the model only on the data stream that is to be compressed (Bellard, 2019; Goyal et al., 2020; Bellard, 2021; Mao et al., 2022). Finally, concurrent work (Valmeekam et al., 2023) also investigated lossless offline compression with foundation models, using arithmetic coding with LLaMA-7B (Touvron et al., 2023).
@@
Much effort has been devoted on understanding the inductive biases of neural networks. Here, we are mostly interested in the biases of Natural Language Processing (NLP) and Transformers. Kudo & Richardson (2018) defined a tokenizer for NLP-related research, an improvement of well-known techniques like byte-pair encoding (BPE) (Sennrich et al., 2016), BPE dropout (Provilkov et al., 2020), and subword regularization (Kudo, 2018). In this paper, we show how these tokenization techniques act as pre-compressors for the data, and can significantly affect the final compression rates when paired with a neural model. More general studies have been performed on generalization (Neyshabur et al., 2017), which, we argue, is equivalent to the model’s compressive power when accounting parameters code-length. Finally, some work has been done on compressing the neural models’ parameters themselves (Cheng et al., 2017).
@@ 5 Conclusion
In this paper we investigated how and why compression and prediction are equivalent. Arithmetic coding transforms a prediction model into a compressor, and, conversely, a compressor can be transformed into a predictor by using the coding lengths to construct probability distributions following Shannon’s entropy principle. We evaluated large pretrained models used as compressors against various standard compressors, and showed they are competitive not only on text but also on modalities they have never been trained on (images, audio data). We showed that the compression viewpoint provides novel insights on scaling laws since it takes the model size into account, unlike the log-loss objective, which is standard in current language modeling research. Consequently, we showed that the optimal model size is inextricably linked to the dataset size and cannot be scaled without limit.
@@ Acknowledgments
We thank Jörg Bornschein, Nando de Freitas, Slav Petrov, and Zhengdong Wang for their helpful feedback and insightful discussions.

### v2

@@ 3.5 Sequential Evolution of In-Context Compression
Language models take a very different “approach” to compression compared to classical compressors. Classical compressors have a small program size and optimize for a large context length to exploit sequential dependencies in the data. In contrast, foundation models consist of billions of parameters, which enable rapid adaptation in their (relatively) short context window (Genewein et al., 2023). Thus, arithmetic coding-based compressors rely heavily on the predictive models’ in-context learning capabilities to achieve competitive compression performance. We investigate this phenomenon in Fig. 4, which visualizes the compression rate across sequence lengths for gzip, Chinchilla 1B and a Transformer pretrained on enwik8. Intuitively, the longer the sequence, the more data the model can process in its context, and therefore, the better the compression. As expected, most compression rates decrease quickly with increasing sequence length, indicating that the models learn some data statistics in-context, without any gradient-based training. As in Table 1, the Chinchilla model achieves the best compression rates across all three data modalities and sequence lengths.
@@ 3.6 Tokenization Is Compression
Increasing the number of tokens (i.e., the “alphabet size”) reduces the length of the sequence and thus increases the amount of information in a models context. However, decreasing the sequence length comes at a price: the number of tokens is larger, which makes the prediction task more challenging since reducing the entropy of the conditional distribution $\rho(x_{i}\mid x_{<i})$ is increasingly difficult for larger alphabet size. In theory, as the tokenization is a lossless compression, the two effects should compensate. In practice, we observe that if the model is small, increasing the number of possible tokens boosts the compression performance. In contrast, for bigger models, it seems that the converse happens: having a larger token vocabulary harms the final compression rate of the model. Nevertheless, short sequence lengths also help Transformers since their time complexity scales quadratically with context length, and it has been shown they do not generalize well to long contexts (Delétang et al., 2023; Ruoss et al., 2023).
@@ 4 Related work
Leveraging Shannon’s source coding theorem (Shannon, 1948), a plethora of approaches exploit the connection between prediction and compression. For example, context-tree weighting (CTW) (Willems et al., 1995) mixes the predictions of many underlying Markov models to achieve lossless compression via arithmetic coding (Pasco, 1977; Rissanen, 1976). Similarly, prediction by partial matching (PPM) (Cleary & Witten, 1984) also leverages arithmetic coding, but uses a contiguous context matching method to create probability distributions based on the history of characters in a sequence. Likewise, PAQ8 (Knoll & de Freitas, 2012) uses a weighted combination of predictions from a large number of models (most of them based on context matching, but unlike PPM also noncontiguous context matches). In a different setting, Veness et al. (2015) demonstrated how to employ compression to obtain value estimates of a policy in an environment. Frank et al. (2000) and later Teahan & Harper (2003) introduced the idea of classification with compressors. Recently, Jiang et al. (2023) applied this technique with NLP tasks, paired with a k-nearest-neighbour algorithm, in which gzip achieves good results. Jiang et al. (2022) exploit the same idea but train the compressor on a vast amount of unlabeled data first. Further comparisons on the same tasks have been done between gzip and simple bag-of-words models (Opitz, 2023). Finally, van den Oord & Schrauwen (2014) apply arithmetic coding to image compression using Student distribution mixtures and Gaussian processes as predictors.
@@
Prior work demonstrated that neural predictive distributions can be employed to perform lossless compression via arithmetic coding (Schmidhuber & Heil, 1994; 1996; Mahoney, 2000; Knoll, 2014; Cox, 2016; Schiopu et al., 2018; Goyal et al., 2019; Liu et al., 2019; Mentzer et al., 2019; 2020; Schiopu & Munteanu, 2020; Rhee et al., 2022; Mikolov, 2012). Similarly, neural networks were also shown to achieve strong lossless compression rates when replacing arithmetic coding with asymmetric numeral systems (Hoogeboom et al., 2019; Kingma et al., 2019; Townsend et al., 2019; Barzen et al., 2022). While these approaches assume the existence of a separate training set, a different line of work investigated arithmetic coding-based neural compression in a purely online fashion, i.e., training the model only on the data stream that is to be compressed (Bellard, 2019; Goyal et al., 2020; Bellard, 2021; Mao et al., 2022). Concurrent work (Valmeekam et al., 2023) also investigated lossless offline compression with foundation models, using arithmetic coding with Llama 2 (7B) (Touvron et al., 2023b).
@@
Much effort has been devoted to understanding the inductive biases of neural networks, in particular with respect to Natural Language Processing (NLP) and Transformers. Kudo & Richardson (2018) developed a tokenizer for NLP that improves over well-known techniques such as byte-pair encoding (BPE) (Sennrich et al., 2016), BPE dropout (Provilkov et al., 2020), and subword regularization (Kudo, 2018). In this paper, we show how these tokenization techniques act as pre-compressors for the data and can significantly affect the final compression rates when paired with a neural model. Other work investigated different aspects of generalization (Neyshabur et al., 2017; Mirchandani et al., 2023; Ge et al., 2023), which is equivalent to compression when accounting for the parameters’ codejlength. Finally, Cheng et al. (2017) investigated compressing the neural models’ parameters to further reduce the code length.
@@ 5 Conclusion
In this paper we investigated how and why sequence modeling is equivalent to compression. Arithmetic coding transforms a sequence model into a compressor, and, conversely, a compressor can be transformed into a predictor using its coding lengths to construct probability distributions following Shannon’s entropy principle. We evaluated large language models as compressors against various standard compressors and showed that they are competitive not only on text but also on modalities they have never been trained on (image and audio data). We showed that the compression view provides novel insights on scaling laws since it takes the model size into account, unlike the log-loss objective, which is standard in current language modeling research. Consequently, we showed that the optimal model size is inextricably linked to the dataset size and cannot be scaled without limit.
@@ Acknowledgments
We thank Jörg Bornschein, Nando de Freitas, Slav Petrov, Zhengdong Wang, and the anonymous reviewers for their helpful feedback and insightful discussions.
Appendix A Arithmetic Coding
Here we provide a step-by-step explanation of the arithmetic encoding example visualized in Fig. 1. Recall from Section 2 that arithmetic encoding iteratively partitions the interval $I=[0,1)$ according to a predictive model $P$ and an input string, i.e., $AIXI$ for Fig. 1.
First, we construct the intervals for the first token, corresponding to $P(\cdot)$:
  • $[0,0.45)$ for $P(A)=0.45$
  • $[0.45,0.75)$ for $P(I)=0.3$
  • $[0.75,1)$ for $P(X)=0.25$
Since the first token is $A$, we set $I=[0,0.45)$ and iterate. Thus, the intervals for $P(\cdot|A)$ are:
  • $[0.2*0,0.2*0.45)=[0,0.09)$ for $P(A|A)=0.2$
  • $[0.09,(0.2+0.6)*0.45)=[0.09,0.36)$ for $P(I|A)=0.3$
  • $[0.36,(0.2+0.6+0.2)*0.45)=[0.36,0.45)$ for $P(X|A)=0.2$
Since the next token is $I$, we set $I=[0,0.45)$ and so on. We terminate with $I=[0.322,0341)$ for $AIXI$. Next, arithmetic will compute the binary sequence corresponding to iteratively splitting the interval $[0,1)$ in half until it is fully contained in $I$. Concretely, this yields the binary sequence.
  • $b0\rightarrow[0,0.5)$
  • $b01\rightarrow[0.25,0.5)$
  • $b010\rightarrow[0.25,0.375)$
  • $b0101\rightarrow[0.3125,0.375)$
  • $b01010\rightarrow[0.3125,0.34375)$
  • $b010101\rightarrow[0.328125,0.34375)$
  • $b0101010\rightarrow[0.328125,0.3359375)$
As $[0.328125,0.3359375)$ is fully contained in $I=[0.322,0341)$, the compressed output is $0101010$, which consists of 7 bits as opposed to the 4 bytes used to encode $AIXI$.
Appendix B Experimental Details
B.1 Data manipulation
ASCII as a standard
The neural models we use (Transformers trained specifically on enwik, and Chinchilla which is trained on a vast and diverse set of text from various sources) take text as input. In theory, they could take in any Unicode character, but unfortunately not all bytes in the range [128, 256]. However, any ASCII character, i.e. in the range [0, 127] is a valid input. Therefore, we chose to compress all bytes sequences by first mapping them to the range [0, 127]. We account for the loss of 1 bit by updating the compression ratio accordingly, if necessary. Each character that needs some processing adds 1 bit to the final compressed sequence.
Compressing text
Text compression is very straightforward. As state above, the text is encoded in ASCII format. If the character is special, we zero the most significant bit, such that the value falls back in the range [0, 127], suitable for ASCII encoding. We append this lost bit at the end of the compressed byte sequence, to account for it in the compression rate.
Compressing images
Image data comes as patches of size (32, 64), each pixel being grayscaled and therefore encoded on 1 byte. This byte gives the brightness of the pixel, 0 being completely black and 255 being completely white. We then flatten this patch to get a sequence of 2048 bytes. Note that it means we lose some 2 dimension correlation between the pixels: the last pixel of the first line will be right next to the first pixel of the second line. For each byte, to map it into the range [0, 127], we simply divide it by 2, and lose the least significant bit. This is performed via simple byte shifting. Note that we do this transformation for every byte, and not only those in the range [128, 255], for consistency. We append the lost bit to the end of the compressed byte sequence, as explained above.
Compressing sound
Sound data from LibriSpeech is sampled at 16Khz, with a sample size of 2 bytes (int16). We manually reduce the sample size to 1 byte. We then split the raw dataset into chunks of size 2048 bytes each, which correspond to roughly 64 milliseconds of speech. We checked whether reducing the sample rate to have longer sequences had a significant impact on the relative compression powers of the models, and we concluded that it had not. We chose to keep the original rate for our experiments. Exactly as for images, for each byte, to map it into the range [0, 127], we simply divide it by 2, and lose the least significant bit. Note that we do this transformation for every byte, and not only those in the range [128, 255], for consistency. The lost bit is also appended at the end of the compressed byte sequence, as explained above.
B.2 Large language models tokenization
As described in the last subsection, the data fed to the large language models we use (Chinchilla and LLama2) is an ASCII string of exactly 2048 characters. However, the models immediately tokenizes the string using SentencePiece (Kudo & Richardson, 2018). The string is transformed into a sequence of integer tokens between 0 and $T$, $T$ being the vocabulary size (they both use $T=32000$). Note that the length of the sequence has now completely changed, and depends on the input: tokenization is already a form of lossless compression. This sequence is fed into the big pretrained Transformer model, which gives us the conditionals $\hat{\rho}(y|x_{<i})$ for all histories $x_{<i}$ and tokens in the alphabet $y$. Denoting the length of the sequence after tokenization as $l$, we obtain $l*T$ log-probabilities. We can pass them to an arithmetic encoder of vocabulary size $T$, to encode the sequence into bits. This is our final compressed sequence, which size in bytes is compared with the initial size, i.e., 2048 bytes.
In practice, the large models had only access to the top-k next token log-probabilities, for each context. We chose $k=100$, which almost fully recovers the conditional distribution. Arithmetic coding can still be applied as the alphabet size is allowed to change while coding: what matters is that the conditional probabilities in each step sum to 1. Accordingly, we renormalize the top-k log-probabilities.
The Transformer models we trained specifically on enwik do not use any tokenization, except in Section 3.6. The reasoning above also holds, except that our models returned the full distribution over tokens, and not only the top-k.
Appendix C Additional Results
Fig. C.1, Fig. 3 and Fig. C.3 show data autoregressively generated by compressors, one step at a time. Note that for Chinchilla, we generate tokens (which size in bytes can vary) until we reach the length in bytes we desire. Also, note that gzip samples are biased, as explained in Section 3.4: looking 1 step ahead is not sufficient to get good samples, and it’s likely that looking multiple steps ahead would improve the results. However, that’s not the purpose of this paper, and we kept the simplest setup for all our compressors.
Context Text (1948 Bytes)
Ground Truth (100 Bytes)
gzip Samples (100 Bytes)
Chinchilla 1B Samples (100 bytes)
Chinchilla 7B Samples (100 bytes)
Chinchilla 70B Samples (100 bytes)
Figure C.1: Compression-based generation for text data. We condition gzip and Chinchilla on a context text of size 1948 bytes (from enwik9) and then sample 100 bytes ($N$ tokens) autoregressively. Since Chinchilla employs a tokenizer, the sampled sequences will contain $N$ tokens, which do not necessarily decode to 100 bytes. Chinchilla’s predictions are significantly more coherent than gzip’s.
(a) Chinchilla 1b
(b) Chinchilla 7b
(c) Chinchilla 70b
Figure C.2: Compression-based generation for image data, for 3 Chinchilla models with different number of parameters. We condition the models on the first half of every row of the image (250 bytes) and then sample the remaining half (250 bytes) autoregressively.
(a) Original spectrogram
(b) gzip
(c) Chinchilla
Figure C.3: Compression-based generation for audio data. We condition gzip and Chinchilla on the first 1024 bytes of the base sequence (from LibriSpeech) and then sample the remaining 1024 bytes autoregressively. Chinchilla predictions exhibit a typical “loop” pattern of autoregressive generation.